Sol Summary Report – February 17th

Crew 275 Sol Summary Report 17Feb2023

Sol: 5

Summary Title: “Mission Support, we have a problem…”

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

After working quite late in the evening on the EVA planned for this morning, we had to wake up way earlier than planned because of another blackout. Again, the generator failed in its fight against the cold. The latter reached us in the Upper Deck of the Hab, where it was only 13°C/56°F at 5:27 am. It is at that time that we contacted Mission Support. They reacted very efficiently, and power was soon back in the Hab. Still, in the Lower Deck, the temperature was close to freezing. The Commander, Crew Engineer, and Health & Safety Officer joined forces to monitor the situation, as the propane alarm started to ring. We made sure that there was no actual gas leak using portable devices found in the RAM. To help the main heating system, we then tried to turn ON the propane heater of the Lower Deck. However, only after about 20 minutes we could finally hear that some propane was reaching the heater. With the propane heater running, the temperature of the Hab became nominal again at around 11:00. Given these events and their consequences (necessity to monitor the situation, lack of sleep, lack of time to prepare the last details of the EVA), we decided to postpone the EVA planned for the morning. The whole crew was relieved about this decision and took this opportunity to get some well-deserved rest, to clean some parts of the station, and to sort the numerous files that are already gathering on our respective computers.

Look Ahead Plan:

With the events of this night / early morning, we decided to not proceed with the EVA planned at 10:00. We would like to postpone it to tomorrow, as the weather seems suitable for this. The other EVA, dedicated to photogrammetry and initially planned for tomorrow, will also be postponed (to next week).

Anomalies in work: Issues with the generator, leading to a blackout during the night. This led to a low temperature (13°C/56°F) in the Upper Deck and the necessity to limit power consumption more than usual during the rest of the day. The dome of the robotic observatory was still open until about noon, but then closed.

Weather: Sunny, low -16°C/3°F, high -1°C/30°F.

Crew Physical Status: Optimal.

EVA: None.

Reports to be filed: Sol Summary, EVA Report, EVA Request, Operations Report, Journalist Report, HSO Report, GreenHab Report, Astronomy Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site